40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10301     Entry time: Wed Jul 30 23:12:05 2014
Author: Koji 
Type: Update 
Category: General 
Subject: IFO patrol 

- The cable for the beat note was disconnected from the frequency counter and reconnected to the spectrum analyzer.

- PMC/IMC had not been locked for 8 hours. 

- PMC was relocked.

- IMC got immediately relocked. Today IMC relocks very fast.
C1:IOO-MC_REFL_OFFSET -0.238
C1:PSL-FSS_INOFFSET -0.94

- Went to the ETMX table. Aligned the oplev beam on the QPD

- The X end green beam was realigned to the cavity.
I can feel that the two mirrors provides quite independent alignment adjustment. VERY NICE.
Green TRX: without PSL Green - 0.612, with PSL green - 0.725

I can clearly see that the mode matching is not ideal. All the higher modes are LG modes!
The input mode is very round.

- Arm cavities were aligned by ASS

- Tested ASX. PZT2 Pitch/Yaw servos run with the previous setting. We still can maximize the transmission by touching PZT1.

- Now Eric joined the activity.

-  Once the beam is aligned what we could lock was LG00/10/20/30.
   We measured the power in LGn0 modes
   LG00: 0.588
   LG10: 0.154
   LG20: 0.053
   LG30: 0.020

   This suggests that the mode-matching ratio is something like 70%

- Q is aligning the PMC. PMC transmission prev 0.783. Basically we could not improve it.
We thought this number can go up to ~0.82 or even ~0.84. We wonder if this comes from the decay of the laser power or reduced visibility?

 

ELOG V3.1.3-